# Context and Diversity Matter: The Emergence of In-Context Learning in World Models

### v1

@@ B.3 Details of Training
All models were trained on NVIDIA A800 GPUs using the AdamW optimizer with the default settings in PyTorch [76].
@@ Appendix C Additional Results
Additional reults of the performances of L2World in Mazes. Table˜5 presents the prediction error (measured as the mean square error between observations) of L2World and the other baselines on the Maze datasets. The prediction errors are fully consistent with the PSNR evaluation; therefore, we report only the PSNR results in the remaining experiments. Table˜6 details the K-step prediction performance, corresponding to the histograms in Figure˜3.
@@
L2World (Maze-32K-L) 2.09e-02 8.01e-03 4.89e-03 3.43e-03 3.13e-03 2.30e-02 7.51e-03 4.82e-03 3.42e-03 3.42e-03
L2World (Maze-32K-S) 1.39e-02 1.18e-02 1.08e-02 9.53e-03 8.96e-03 1.43e-02 1.19e-02 1.09e-02 9.35e-03 9.31e-03
L2World (Maze-128-S) 1.13e-02 9.15e-03 8.74e-03 6.28e-03 6.66e-03 1.58e-02 1.37e-02 1.26e-02 1.08e-02 1.09e-02
L2World (Maze-128-L) 1.40e-02 8.20e-03 4.66e-03 2.72e-03 2.51e-03 1.76e-02 1.14e-02 8.02e-03 7.01e-03 7.04e-03
Dreamer (Maze-32K-L) 2.29e-02 6.57e-03 1.19e-02 7.47e-03 6.47e-03 2.08e-02 8.95e-03 7.25e-03 5.43e-03 6.14e-03
Dreamer (Maze-128-L) 1.94e-02 8.63e-03 6.56e-03 6.02e-03 5.72e-03 3.75e-02 3.51e-02 3.90e-02 4.50e-02 4.46e-02
NWM (Maze-32K-L) 8.22e-03 9.52e-03 1.20e-02 5.86e-03 7.83e-03 2.02e-02 2.26e-02 2.48e-02 1.70e-02 1.28e-02
@@
ICEL and ICER in navigation world models implicitly perform global mapping. We further show that predicting transitions alone, without any specifically designed tasks, can potentially capture the global map implicitly. We collect $12$ trajectories from $4$ unseen mazes ($3$ trajectories in each maze) and track the transformation of memory states ($\phi_{t}$) across each of the linear attention layers. Among the $18$ layers used in the transition model, we select layers $\{1,6,12,18\}$ for t-SNE visualization of the memory states ($\phi_{t}$). To ensure that the $4$ unseen mazes are not trivially discriminable, for instance, by geometric and embodiment configurations in one-shot, we maintain identical configurations for most aspects of the evaluated $4$ mazes while varying only their topology, such as the arrangement of walls. We employ silhouette scores [77] to quantify clustering quality, where higher values indicate better environment separation. As visualized in Figure˜10, we highlight two insights: first, training solely on transition prediction across diverse environments and long contextual windows implicitly learns a spatial map, potentially removing the need for auxiliary mapping modules; second, ICEL and ICER behave differently across layers, suggesting distinct underlying mechanisms.
@@
Additional Cases. Figure˜11 shows two failure predictions of L2World (Maze-32K-L) on unseen maze environments. These failures appear to stem from excessive blurring in the predictions, which causes the compound error to explode rapidly as $k$ increases. Building on this observation, a natural extension of our work is to inject additional overshooting during training so that the world model can better forecast distant futures. Figure˜12 juxtaposes the performance of L2World trained with 32K-L and 32K-S, clearly evidencing the benefit of longer sequences in incentivizing ICEL.

### v2

@@ B.3 Details of Training
All models were trained on NVIDIA A800 GPUs using the AdamW optimizer with the default settings in PyTorch Paszke et al. (2019).
@@ Appendix C Additional Results
Additional reults of the performances of L2World in Mazes. Table˜5 presents the prediction error (measured as the mean square error between observations) of L2World and the other baselines on the Maze datasets. The prediction errors are fully consistent with the PSNR evaluation; therefore, we report only the PSNR results in the remaining experiments. Table˜6 details the K-step prediction performance, corresponding to the histograms in Figure˜4.
@@
L2World (Maze-32K-L) $2.09\times 10^{-2}$ $8.01\times 10^{-3}$ $4.89\times 10^{-3}$ $3.43\times 10^{-3}$ $3.13\times 10^{-3}$ $2.30\times 10^{-2}$ $\bf{7.51\times 10^{-3}}$ $\bf{4.82\times 10^{-3}}$ $\bf{3.42\times 10^{-3}}$ $\bf{3.42\times 10^{-3}}$
L2World (Maze-32K-S) $39\times 10^{-2} 1.1$$18\times 10^{-2}$ $1.08\times 10^{-2}$ $9.53\times 10^{-3}$ $8.96\times 10^{-3}$ $\bf{1.43\times 10^{-2}}$ $1.19\times 10^{-2}$ $1.09\times 10^{-2}$ $9.35\times 10^{-3}$ $9.31\times 10^{-3}$
L2World (Maze-128-S) $1.13\times 10^{-2}$ $9.15\times 10^{-3}$ $8.74\times 10^{-3}$ $28\times 10^{-3} 6.6$$66\times 10^{-3}$ $1.58\times 10^{-2}$ $1.37\times 10^{-2}$ $1.26\times 10^{-2}$ $1.08\times 10^{-2}$ $1.09\times 10^{-2}$
L2World (Maze-128-L) $1.40\times 10^{-2}$ $8.20\times 10^{-3}$ $\bf{4.66\times 10^{-3}}$ $\bf{2.72\times 10^{-3}}$ $\bf{2.51\times 10^{-3}}$ $76\times 10^{-2} 1.1$$14\times 10^{-2}$ $8.02\times 10^{-3}$ $7.01\times 10^{-3}$ $7.04\times 10^{-3}$
Dreamer (Maze-32K-L) $2.29\times 10^{-2}$ $\bf{6.57\times 10^{-3}}$ $1.19\times 10^{-2}$ $7.47\times 10^{-3}$ $6.47\times 10^{-3}$ $2.08\times 10^{-2}$ $8.95\times 10^{-3}$ $7.25\times 10^{-3}$ $5.43\times 10^{-3}$ $6.14\times 10^{-3}$
Dreamer (Maze-128-L) $1.94\times 10^{-2}$ $8.63\times 10^{-3}$ $6.56\times 10^{-3}$ $6.02\times 10^{-3}$ $5.72\times 10^{-3}$ $3.75\times 10^{-2}$ $3.51\times 10^{-2}$ $3.90\times 10^{-2}$ $50\times 10^{-2} 4.4$$46\times 10^{-2}$
NWM (Maze-32K-L) $\bf{8.22\times 10^{-3}}$ $9.52\times 10^{-3}$ $1.20\times 10^{-2}$ $5.86\times 10^{-3}$ $7.83\times 10^{-3}$ $02\times 10^{-2} 2.2$$26\times 10^{-2}$ $2.48\times 10^{-2}$ $1.70\times 10^{-2}$ $1.28\times 10^{-2}$
@@
EL and ER in navigation world models implicitly perform global mapping. We further show that predicting transitions alone, without any specifically designed tasks, can potentially capture the global map implicitly. We collect $12$ trajectories from $4$ unseen mazes ($3$ trajectories in each maze) and track the transformation of memory states ($\phi_{t}$) across each of the linear attention layers. Among the $18$ layers used in the transition model, we select layers $\{1,6,12,18\}$ for t-SNE visualization of the memory states ($\phi_{t}$). To ensure that the $4$ unseen mazes are not trivially discriminable, for instance, by geometric and embodiment configurations in one-shot, we maintain identical configurations for most aspects of the evaluated $4$ mazes while varying only their topology, such as the arrangement of walls. We employ silhouette scores Rousseeuw (1987) to quantify clustering quality, where higher values indicate better environment separation. As visualized in Figure˜11, we highlight two insights: first, training solely on transition prediction across diverse environments and long contextual windows implicitly learns a spatial map, potentially removing the need for auxiliary mapping modules; second, EL and ER behave differently across layers, suggesting distinct underlying mechanisms.
@@
Cases. Figure˜12 illustrates 10-step-ahead predictions at $T=\{1,100,10K\}$ for our method, Dreamer, and NWM, all trained on Maze-32K-L. NWM produces visually convincing frames with fine textures and a plausible layout; however, frame-wise fidelity alone is insufficient for accurate long-range forecasting because the model lacks long-term memory and spatial reasoning. Figure˜13 juxtaposes the performance of L2World trained with 32K-L and 32K-S, clearly demonstrating the benefit of longer sequences in promoting EL. Figure˜14 shows two failure cases of L2World (Maze-32K-L) on unseen maze environments. These failures appear to result from excessive blurring in the predictions, causing the compound error to escalate rapidly as $k$ increases. Building on this observation, a natural extension of our work is to incorporate additional overshooting during training so that the world model can better forecast distant futures.
Appendix D Use of LLMs
We used large language models (LLMs) only as an auxiliary tool to improve the clarity and presentation of this paper. The assistance was limited to:
  • Language refinement: grammar checking, wording suggestions, and improving sentence fluency while preserving the authors’ original technical content.
  • Mathematical support: helping verify the correctness and readability of some derivations and notations, without introducing new technical results.
No LLM was used for generating research ideas, designing experiments, analyzing results, or writing original scientific content. All conceptual and technical contributions were made by the authors.